# Neural variational inference for cutting feedback during uncertainty propagation

### v1

@@ Appendix B Proofs
Appendix C Proofs for neural flow examples
We now prove the results for specific neural flow classes.
C.1 Common network notation and quantitative bounds
Let $K_{M}=K_{\eta}\times[-M,M]$, $d_{\eta}:=\dim(\operatorname{aff}(K_{\eta}))$, dimension of the affine hull of the set $K_{\eta}$, and $\operatorname{diam}_{\eta}:=\sup_{\eta,\eta^{\prime}\in K_{\eta}}\|\eta-\eta^{\prime}\|_{2}$. All suprema are over $K_{M}$. Write $\|\,\cdot\,\|_{\infty}$ for the max norm. Throughout, clipping uses the one Lipschitz map $\mathrm{clip}(x;a,b):=\min\{b,\max\{a,x\}\}$. Consider a fully connected ReLU network $h_{\vartheta}:\mathbb{R}^{d}\to\mathbb{R}$ with depth (number of layers) $L$ and maximal width $W$. Let $W_{\ell}(\vartheta)$ denote the layer matrices after absorbing biases as an extra column. Set $S(\vartheta):=\max_{\ell}\|W_{\ell}(\vartheta)\|_{\mathrm{op}}$ and $S_{\Pi}(\vartheta):=\prod_{\ell=1}^{L}\|W_{\ell}(\vartheta)\|_{\mathrm{op}}$. Then by chain rule and as derivative of ReLU (where it exists) is less than $1$, we have $\sup_{x}\|\nabla h_{\vartheta}(x)\|\leq S_{\Pi}(\vartheta)\leq S(\vartheta)^{L}$. If $\vartheta,\vartheta^{\prime}$ lie in the same axis aligned box $B,B]^{m}$, then each layer perturbation satisfies $\|W_{\ell}(\vartheta)-W_{\ell}(\vartheta^{\prime})\|_{\mathrm{op}}\leq RW\|\vartheta-\vartheta^{\prime}\|_{\infty}$. A layerwise chain rule with ReLU one Lipschitz gives, for every input ball $\|x\|\leq B$,
$\sup_{\|x\|\leq B}\big|h_{\vartheta}(x)-h_{\vartheta^{\prime}}(x)\big|\;\leq\;L\,W\,B\,\big(RW\big)^{L}\,\|\vartheta-\vartheta^{\prime}\|_{\infty}.$
C.2 Unconstrained Monotonic Neural Networks
@@
It is enough to show that the class satisfy Assumption 2 as the rest follows from Theorem 2. Define
$T_{\vartheta}(\eta,z)$ $=a_{\vartheta}(\eta)+\int_{0}^{z}\exp\!\big(g_{\vartheta}(\eta,t)\big)\,dt,$
$\qquad a_{\vartheta}$ $=\mathrm{clip}(\tilde{a}_{\vartheta},-M^{*},M^{*}),$
$\qquad g_{\vartheta}$ $=\mathrm{clip}(\tilde{g}_{\vartheta},-A^{*}-B^{*}|z|,A^{*}+B^{*}|z|),$
where $\tilde{a}_{\vartheta}$ is a ReLU network on $K_{\eta}$ and $\tilde{g}_{\vartheta}$ is a ReLU network on $K_{M}$. The clips are pointwise in the displayed variables. Since clipping is one Lipschitz, it preserves continuity and does not increase moduli.
For every $(\eta,z)$ one has $\partial_{z}T_{\vartheta}(\eta,z)=\exp\big(g_{\vartheta}(\eta,z)\big)>0$. Continuity in $(\eta,z)$ follows from continuity of $\tilde{a}_{\vartheta}$ and $\tilde{g}_{\vartheta}$ and of the exponential and of the integral in the upper limit. Differentiability in $z$ and continuity of this partial derivative follows directly from the fundamental theorem of calculus, yielding $\partial_{z}T_{\vartheta}(\eta,z)=\exp(g_{\vartheta}(\eta,z))$. Finally, compactness of the parameter set is ensured by clipping the weights to [B,B][-B,B] giving $\Theta_{m}=[-B,B]^{m}$. So, Assumption 2(i) is satisfied.
By definition $|T_{\vartheta}(\eta,0)|=|a_{\vartheta}(\eta)|\leq M^{*}$. Also $|\log\partial_{z}T_{\vartheta}(\eta,z)|=|g_{\vartheta}(\eta,z)|\leq A^{*}+B^{*}|z|$. This is precisely the global linear envelope. So, Assumption 2(ii) is satisfied.
Lastly, we establish part (iii) with an explicit construction. Let the target quantile map $T_{p}$ satisfy Definition 1 on $K_{M}$. Define $u_{1}(\eta):=T_{p}(\eta,0)$ and $u_{2}(\eta,z):=\log\partial_{z}T_{p}(\eta,z)$. The definition yields constants $M_{0}>0$ and $A_{d},B_{d},C_{d}\geq 0$ with
$\sup_{\eta\in K_{\eta}}|u_{1}(\eta)|\leq M_{0},\qquad\sup_{(\eta,z)\in K_{M}}|u_{2}(\eta,z)|\leq A_{d}+B_{d}|z|.$
Choose thresholds $M^{*}>M_{0}$, $A^{*}>A_{d}$, $B^{*}>B_{d}$. Hence if we construct $\tilde{a}_{\vartheta},\tilde{g}_{\vartheta}$ so that
$\sup_{\eta\in K_{\eta}}\big|\tilde{a}_{\vartheta}(\eta)-u_{1}(\eta)\big|\leq\varepsilon,\qquad\sup_{(\eta,z)\in K_{M}}\big|\tilde{g}_{\vartheta}(\eta,z)-u_{2}(\eta,z)\big|\leq\varepsilon,\qquad 0<\varepsilon\leq\varepsilon_{\max},$
with $\varepsilon_{\max}:=\min\{M^{*}-M_{0},\ A^{*}-A_{d},\ B^{*}-B_{d}\}$, then the clips are inactive on $K_{M}$ and $a_{\vartheta}=\tilde{a}_{\vartheta}$, $g_{\vartheta}=\tilde{g}_{\vartheta}$ there. It follows that
$\sup_{\eta\in K_{\eta}}|a_{\vartheta}(\eta)-u_{1}(\eta)|\leq\varepsilon,\qquad\sup_{(\eta,z)\in K_{M}}|g_{\vartheta}(\eta,z)-u_{2}(\eta,z)|\leq\varepsilon.$
To realize $(\mathrm{Apx})$ with quantitative size, note that Definition 1 supplies a Lipschitz moduli $\L _{a}$ for $u_{1}$ on $K_{\eta}$ and $\L _{b}(M)$ for $u_{2}$. We approximate the true maps
@@
with quantitative budgets that depend on $L_{a}$ and $L_{b}(M)$.
Theorem 1 in Yarotsky, (2017) proves that for any $d\geq 1$, a Lipschitz function $f$ on $0,1]^{d}$ with $\|f\|_{\infty}\leq 1$ admits a ReLU network approximation with depth $O(-\log(\varepsilon))$ and number of weights $O\Big(\varepsilon^{-d}[-\log(\varepsilon)]\Big)$ that achieves sup norm error at most $\varepsilon$ on $0,1]^{d}$. As $u_{1}$ is $L_{a}$-Lipschitz on $K_{\eta}$ and $u_{2}$ is $L_{b}(M)$-Lipschitz on $K_{M}$, we map each domain to the unit cube by the affine scaling $x\mapsto(x-x_{0})/\operatorname{diam}$. Lipschitz constants scale by the domain diameter, so the rescaled functions are $L_{a}\,\operatorname{diam}_{\eta}$ and $L_{b}(M)(\operatorname{diam}_{\eta}+M)$ Lipschitz on $0,1]^{d}$ up to absolute constants. Applying the theorem of Yarotsky, (2017) and scaling the error back gives realizations $\tilde{a}_{\vartheta}$ and $\tilde{g}_{\vartheta}$ with
$\|\tilde{a}_{\vartheta}-u_{1}\|_{\infty,K_{\eta}}\leq\varepsilon,\qquad\|\tilde{g}_{\vartheta}-u_{2}\|_{\infty,K_{M}}\leq\varepsilon,$
@@
$\tilde{a}_{\vartheta}\;\leq\;C_{a}\Bigl(\tfrac{L_{a}\,\operatorname{diam}_{\eta}}{\varepsilon}\Bigr)^{d_{\eta}}\log\tfrac{1}{\varepsilon},\qquad\#\tilde{g}_{\vartheta}\;\leq\;C_{g}\Bigl(\tfrac{L_{b}(M)(\operatorname{diam}_{\eta}+M)}{\varepsilon}\Bigr)^{d_{\eta}+1}\log\tfrac{1}{\varepsilon}.$
Here $C_{a}$ and $C_{g}$ are fixed constants depending only on $d_{\eta}$. This completes the verification for UMNN. ∎
C.3 Rational Quadratic Neural Spline Flows
@@
Partition [M,M][-M,M] into $2K$ equal bins of width $w=M/K$ with knots $z_{i}=-M+jw$ for $j=0,\ldots,2K$, so $z_{K}=0$. We will choose $K$ later and maintain $K\geq M$, hence $w\leq 1$. For each $\eta$ the conditioner $\psi_{\vartheta}(\eta)$ outputs raw vectors $b^{(h)}(\eta)\in\mathbb{R}^{2K}$, $v(\eta)\in\mathbb{R}^{2K+1}$, and a scalar offset $c(\eta)$ with $|c(\eta)|\leq B_{0}$. Define heights with a total mass $H_{\mathrm{tot}}>0$,
$h_{i}(\eta)=H_{\mathrm{tot}}\ \mathrm{softmax}\!\big(b^{(h)}(\eta)\big)_{i},\qquad y_{0}:=b(\eta),\quad y_{i+1}:=y_{i}+h_{i}(\eta)\in[b(\eta),\,b(\eta)+H_{\mathrm{tot}}],$
@@
The numerator of $\partial_{z}T_{\vartheta}(\eta,z)$ is strictly positive for $t\in[0,1]$ since $s_{i},s_{i+1}>0$, so the derivative is positive on each bin. By construction of RQ-NSF, for adjacent bins, the value and the first derivative match at the knots. Extending with linear tails $z\mapsto y_{0}+s_{0}(z-z_{0})$ for $z\leq z_{0}$ and $z\mapsto y_{2K}+s_{2K}(z-z_{2K})$ for $z\geq z_{2K}$ yields a globally $C^{1}$ strictly increasing map. This proves Assumption 2(i).
@@
$|T_{\vartheta}(\eta,0)|=|y_{K}(\eta)|\leq|b(\eta)|+\sum_{i=0}^{2K-1}h_{i}(\eta)\leq B_{0}+H_{\mathrm{tot}}=:M^{*}.$
Choosing $B_{0}\geq M_{0}$ establishes Assumption 2(ii).
Let $T_{p}$ satisfy Definition 1 and set $u(\eta,z):=\log\partial_{z}T_{p}(\eta,z)$. By Lemma 1, on $K_{M}=[-M,M]\times K_{\eta}$ the function $u$ is $L_{b}(M)$–Lipschitz with
$L_{b}(M)=M+e^{H(M)}L_{S}(M)+L_{CP}(M).$
Using the RQS formula with endpoint logits $\beta_{i}(\eta)=u(\eta,z_{i})$ on each bin,
$\sup_{\tau\in[0,w]}\big|\log\partial_{z}T_{\vartheta}(\eta,z_{i}+\tau)-u(\eta,z_{i}+\tau)\big|\leq L_{b}(M)\,w,$
@@
$\sup_{(\eta,z)\in K_{M}}\big|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\big|\leq L_{b}(M)\frac{M}{K}.$
Impose the second inequality in Assumption 2(iii) with tolerance $\varepsilon$ by choosing
$K_{\star}:=\Big\lceil\tfrac{L_{b}(M)\,M}{\varepsilon}\Big\rceil,$
@@
Moreover, Lemma 1 gives $\mathrm{Lip}_{\eta}(u(\cdot,z))\leq L_{b}(M)$; hence
$|h_{i}^{\star}(\eta)-h_{i}^{\star}(\eta^{\prime})|$ $\leq\int_{z_{i}}^{z_{i+1}}e^{H(M)}\,|u(\eta,t)-u(\eta^{\prime},t)|\,dt\leq w\,e^{H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|,$
$|\log h_{i}^{\star}(\eta)-\log h_{i}^{\star}(\eta^{\prime})|$ $\leq\frac{|h_{i}^{\star}(\eta)-h_{i}^{\star}(\eta^{\prime})|}{\min h_{i}^{\star}}\leq e^{2H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|.$
We have $2Me^{-H(M)}\leq S(\eta)\leq 2Me^{H(M)}$ and $\mathrm{Lip}_{\eta}(S)\leq 2Me^{H(M)}L_{b}(M)$, hence
$|a(\eta)-a(\eta^{\prime})|=|\log S(\eta)-\log S(\eta^{\prime})|\leq e^{2H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|.$
@@
$|b_{i}^{\star}(\eta)-b_{i}^{\star}(\eta^{\prime})|\leq 2\,e^{2H(M)}L_{b}(M)\,\|\eta-\eta^{\prime}\|\quad\Rightarrow\quad\mathrm{Lip}_{\eta}\big(b^{\star}\big)\leq 2\,e^{2H(M)}L_{b}(M).$
Set
$F(\eta)=\big(b^{\star}(\eta),\,v^{\star}(\eta),\,c^{\star}(\eta)\big)\in\mathbb{R}^{4K_{\star}+2},\qquad v_{i}^{\star}(\eta)=u(\eta,z_{i}),\quad c^{\star}(\eta)=T_{p}(\eta,-M).$
So, $b^{\star}$ is $2e^{2H(M)}L_{b}(M)$–Lipschitz and by Lemma 1 $v^{\star}$ is $L_{b}(M)$–Lipschitz. We now establish the Lipschitz constant for $c^{*}$. We start from the quantile identity
@@
By Definition 1(ii), for $|\theta|\leq G(M)$ we have $\|\nabla_{\eta}\log p(\theta\mid\eta)\|\leq L_{CP}(G(M))$. Quantile growth ensures $|T_{p}(\eta,z)|\leq G(|z|)\leq G(M)$ for all $z\in(-\infty,-M]$. Therefore
@@
From Assumption 2(ii), $|\log\partial_{z}T_{p}(\eta,z)|\leq H(|z|)$, hence $\partial_{z}T_{p}(\eta,-M)\leq e^{H(M)}$, so
@@
Combining and applying Mill’s ratio, $c^{\star}$ is $L_{\mathrm{edge}}(M)$–Lipschitz with
@@
By the ReLU vector sup–norm approximation theorem (Yarotsky,, 2017), applied to $F=(b^{\star},v^{\star},c^{\star})$ with
$\mathrm{Lip}_{\eta}(F)\ \leq\ 2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\qquad\text{where}\qquad L_{\mathrm{edge}}(M)=\frac{e^{H(M)}}{M}\,L_{CP}\!\big(G(M)\big),$
there exists a conditioner $\psi_{\vartheta}:K_{\eta}\to\mathbb{R}^{4K_{\star}+2}$ such that
@@
$m\ \leq\ C_{r}\,(4K_{\star}+2)\,A^{d_{\eta}}\log A,\qquad A\ :=\ \frac{\big(2e^{2H(M)}L_{b}(M)+L_{\mathrm{edge}}(M)\big)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}.$
@@
$\delta\coloneqq\sup_{(\eta,z)\in K_{M}}\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|.$
@@
$\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|\leq L_{b}(M)\,w+\tau,\qquad w=\frac{M}{K_{\star}},$
@@
$K_{\star}=\Bigl\lceil\frac{L_{b}(M)M}{\varepsilon_{d}}\Bigr\rceil\qquad\text{with}\qquad\varepsilon_{d}=\frac{\varepsilon}{4\,M\,e^{H(M)}},$
so that $L_{b}(M)w\leq\varepsilon_{d}$. Then
$\delta\ \leq\ L_{b}(M)w+\tau\ \leq\ \varepsilon_{d}+\varepsilon_{\mathrm{par}}\ \leq\ \frac{\varepsilon}{4Me^{H(M)}}+\frac{\varepsilon}{4Me^{H(M)}}\ =\ \frac{\varepsilon}{2Me^{H(M)}},$
@@
$\sup_{(\eta,z)\in K_{M}}\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|\ \leq\ \varepsilon.$
@@
Using $\delta\leq\varepsilon/(2Me^{H(M)})$ and $e^{x}-1\leq x$ for $x\geq 0$ gives $Me^{H(M)}(e^{\delta}-1)\leq\varepsilon/2$, and since $\varepsilon_{\mathrm{par}}\leq\varepsilon/2$,
@@
Thus both inequalities in Assumption 2(iii) hold with tolerance $\varepsilon$.

### v2

@@ Acknowledgements
Appendix S5 Proofs for Neural Flow Examples
We now prove Theorems 2 and 3 for specific neural flow classes.
S5.1 Common Network Notation and Quantitative Bounds
Let $\eta\in$ some compact $K_{\eta}\in\mathbb{R}^{d}$ and $\operatorname{diam}_{\eta}:=\sup_{\eta,\eta^{\prime}\in K_{\eta}}\|\eta-\eta^{\prime}\|_{2}$. Write $\|\,\cdot\,\|_{\infty}$ for the max norm and $\|\cdot\|_{\text{op}}$ for the operator norm. Throughout, clipping uses the one Lipschitz map $\mathrm{clip}(x;a,b):=\min\{b,\max\{a,x\}\}$.
S5.2 Unconstrained Monotonic Neural Networks
@@
It is enough to show that the class satisfies Assumption 2 as the rest follows from Theorem 2. Define
$\begin{split}T_{\vartheta}(\eta,z)&=a_{\vartheta_{a}}(\eta)+\int_{0}^{z}\exp\!\big(g_{\vartheta_{g}}(\eta,t)\big)\,dt,\\ \qquad a_{\vartheta_{a}}&=\mathrm{clip}(\tilde{a}_{\vartheta_{a}},-M^{*},M^{*}),\\ \qquad g_{\vartheta_{g}}&=\mathrm{clip}(\tilde{g}_{\vartheta_{g}},-A^{*}-B^{*}|z|,A^{*}+B^{*}|z|),\end{split}$ (S10)
where $\tilde{a}_{\vartheta_{a}}$ is a ReLU network on $K_{\eta}$, $\tilde{g}_{\vartheta_{g}}$ is a ReLU network on $K_{M}$, and $\vartheta=(\vartheta_{a},\vartheta_{g})^{\top}$. The clips are pointwise in the displayed variables. Since clipping is one Lipschitz, it preserves continuity and does not increase moduli.
Assumption 2(i)(i).
For every $(\eta,z,\vartheta)$ one has $\partial_{z}T_{\vartheta}(\eta,z)=\exp\big(g_{\vartheta_{g}}(\eta,z)\big)>0$. Continuity in $(\eta,z,\vartheta)$ follows from continuity of $\tilde{a}_{\vartheta_{a}}$, $\tilde{g}_{\vartheta_{g}}$, exponential, and the integral in the upper limit. Differentiability in $z$ and continuity of this partial derivative follow directly from the fundamental theorem of calculus, yielding $\partial_{z}T_{\vartheta}(\eta,z)=\exp(g_{\vartheta_{g}}(\eta,z))$.
Assumption 2(ii)(ii).
Following definition (S10), $|T_{\vartheta}(\eta,0)|=|a_{\vartheta_{a}}(\eta)|\leq M^{*}$ and $|\log\partial_{z}T_{\vartheta}(\eta,z)|=|g_{\vartheta_{g}}(\eta,z)|\leq A^{*}+B^{*}|z|$. This is precisely the global linear envelope. So, Assumption 2(ii)(ii) is satisfied.
Assumption 2(iii)(iii).
Lastly, we establish part $(iii)$ with an explicit construction. Let the target quantile map $T_{p}$ satisfy Assumption 1 on $K_{M}$. Define $u_{1}(\eta):=T_{p}(\eta,0)$ and $u_{2}(\eta,z):=\log\partial_{z}T_{p}(\eta,z)$. The definition yields constants $M_{0}>0$ and $A_{d},B_{d},C_{d}\geq 0$ with
$\sup_{\eta\in K_{\eta}}|u_{1}(\eta)|\leq M_{0},\qquad\sup_{(z,\eta)\in K_{M}}|u_{2}(\eta,z)|\leq A_{d}+B_{d}|z|.$
Choose thresholds $M^{*}>M_{0}$, $A^{*}>A_{d}$, $B^{*}>B_{d}$. Hence if we construct $\tilde{a}_{\vartheta_{a}},\tilde{g}_{\vartheta_{g}}$ so that
$\sup_{\eta\in K_{\eta}}\big|\tilde{a}_{\vartheta_{a}}(\eta)-u_{1}(\eta)\big|\leq\varepsilon,\qquad\sup_{(z,\eta)\in K_{M}}\big|\tilde{g}_{\vartheta_{g}}(\eta,z)-u_{2}(\eta,z)\big|\leq\varepsilon,\qquad 0<\varepsilon\leq\varepsilon_{\max},$
with $\varepsilon_{\max}:=\min\{M^{*}-M_{0},\ A^{*}-A_{d},\ B^{*}-B_{d}\}$, then the clips are inactive on $K_{M}$ and $a_{\vartheta_{a}}=\tilde{a}_{\vartheta_{a}}$, $g_{\vartheta_{g}}=\tilde{g}_{\vartheta_{g}}$ there. It follows that
$\sup_{\eta\in K_{\eta}}|a_{\vartheta_{a}}(\eta)-u_{1}(\eta)|\leq\varepsilon,\qquad\sup_{(z,\eta)\in K_{M}}|g_{\vartheta}(\eta,z)-u_{2}(\eta,z)|\leq\varepsilon.$
Assumption 1 supplies a Lipschitz moduli $\mathsf{L}_{a}$ for $u_{1}$ on $K_{\eta}$ and $\mathsf{L}_{b}(M)$ for $u_{2}$ on $K_{M}$. That is, we approximate the true maps
@@
with quantitative budgets that depend on $\mathsf{L}_{a}$ and $\mathsf{L}_{b}(M)$.
Theorem 1 in Yarotsky, (2017) proves that for any $d\geq 1$, a 1-Lipschitz function $f$ on $0,1]^{d}$ with $\|f\|_{\infty}\leq 1$ admits a ReLU network approximation with depth $O\left(\log\left(\frac{1}{\varepsilon}\right)\right)$ and number of weights $O\left(\varepsilon^{-d}\log\left(\frac{1}{\varepsilon}\right)\right)$ that achieves sup norm error at most $\varepsilon$ on $0,1]^{d}$. As $u_{1}$ is $\mathsf{L}_{a}$-Lipschitz on $K_{\eta}$ and $u_{2}$ is $\mathsf{L}_{b}(M)$-Lipschitz on $K_{M}$, we map each domain to the unit cube by the affine scaling $x\mapsto(x-x_{0})/\operatorname{diam}$. Lipschitz constants scale by the domain diameter, so the rescaled functions are $\mathsf{L}_{a}\,\operatorname{diam}_{\eta}$ and $\mathsf{L}_{b}(M)(\operatorname{diam}_{\eta}+M)$ Lipschitz on $0,1]^{d}$ up to absolute constants. Applying the theorem of Yarotsky, (2017) and scaling the error back gives realizations $\tilde{a}_{\vartheta_{a}}$ and $\tilde{g}_{\vartheta_{g}}$ with
$\|\tilde{a}_{\vartheta_{a}}-u_{1}\|_{\infty,K_{\eta}}\leq\varepsilon,\qquad\|\tilde{g}_{\vartheta_{g}}-u_{2}\|_{\infty,K_{M}}\leq\varepsilon,$
@@
$\tilde{a}_{\vartheta_{a}}\;\leq\;C_{a}\Bigl(\tfrac{\mathsf{L}_{a}\,\operatorname{diam}_{\eta}}{\varepsilon}\Bigr)^{d_{\eta}}\log\tfrac{1}{\varepsilon},\qquad\#\tilde{g}_{\vartheta_{g}}\;\leq\;C_{g}\Bigl(\tfrac{\mathsf{L}_{b}(M)(\operatorname{diam}_{\eta}+M)}{\varepsilon}\Bigr)^{d_{\eta}+1}\log\tfrac{1}{\varepsilon}.$
Here $C_{a}$ and $C_{g}$ are fixed constants depending only on $d_{\eta}$. This completes the verification of Assumption 2 for UMNN. Then Theorem 2 is proved by simply using $\varepsilon/O(\tilde{L}_{S}(\tilde{G}(M(\varepsilon)))$ which is the required scaling from Theorem 1. ∎
S5.3 Rational Quadratic Neural Spline Flows
@@
Partition [M,M][-M,M] into $2K$ equal bins of width $w=M/K$ with knots $z_{i}=-M+jw$ for $j=0,\ldots,2K$, so $z_{K}=0$. We will choose $K$ later and maintain $K\geq M$, hence $w\leq 1$. For each $\eta$ the conditioner $\psi_{\vartheta}(\eta)$ outputs raw vectors $b^{(h)}(\eta)\in\mathbb{R}^{2K}$, $v(\eta)\in\mathbb{R}^{2K+1}$, and a scalar offset $c(\eta)$ with $|c(\eta)|\leq$ some $B_{0}$ that we will specify later. Define heights with a total mass $H_{\mathrm{tot}}>0$,
$h_{i}(\eta)=H_{\mathrm{tot}}\ \mathrm{softmax}\big(b^{(h)}(\eta)\big)_{i},\quad y_{0}:=c(\eta),\quad y_{i+1}:=y_{i}+h_{i}(\eta)\in[c(\eta),\,c(\eta)+H_{\mathrm{tot}}],$
@@
Assumption 2(i)(i).
The numerator of $\partial_{z}T_{\vartheta}(\eta,z)$ is strictly positive for $t\in[0,1]$ since $s_{i},s_{i+1}>0$, so the derivative is positive on each bin. By construction of RQ-NSF, for adjacent bins, the value and the first derivative match at the knots. Extending with linear tails $z\mapsto y_{0}+s_{0}(z-z_{0})$ for $z\leq z_{0}$ and $z\mapsto y_{2K}+s_{2K}(z-z_{2K})$ for $z\geq z_{2K}$ yields a globally $C^{1}$ strictly increasing map. This proves Assumption 2(i)(i).
Assumption 2(ii)(ii).
@@
$|T_{\vartheta}(\eta,0)|=|y_{K}(\eta)|\leq|c(\eta)|+\sum_{i=0}^{2K-1}h_{i}(\eta)\leq B_{0}+H_{\mathrm{tot}}=:M^{*}.$
Choosing $B_{0}\geq M_{0}$ establishes Assumption 2(ii)(ii).
Assumption 2(iii)(iii).
Let $T_{p}$ satisfy Assumption 1 and set $u(\eta,z):=\log\partial_{z}T_{p}(\eta,z)$. By Lemma 1, on $K_{M}$ the function $u$ is $\mathsf{L}_{b}(M)$-Lipschitz. Using the RQS formula with endpoint logits $\beta_{i}(\eta)=u(\eta,z_{i})$ on each bin,
$\sup_{\tau\in[0,w]}\big|\log\partial_{z}T_{\vartheta}(\eta,z_{i}+\tau)-u(\eta,z_{i}+\tau)\big|\leq\mathsf{L}_{b}(M)\,w.$
@@
$\sup_{(z,\eta)\in K_{M}}\big|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\big|\leq\mathsf{L}_{b}(M)\frac{M}{K}.$
Impose the second inequality in Assumption 2(iii)(iii) with tolerance $\varepsilon$ by choosing
$K_{\star}:=\Big\lceil\tfrac{\mathsf{L}_{b}(M)\,M}{\varepsilon}\Big\rceil,$
@@
Moreover, Lemma 1 gives $\mathrm{Lip}_{\eta}(u(\cdot,z))\leq\mathsf{L}_{b}(M)$; hence
$|h_{i}^{\star}(\eta)-h_{i}^{\star}(\eta^{\prime})|$ $\leq\int_{z_{i}}^{z_{i+1}}e^{H(M)}\,|u(\eta,t)-u(\eta^{\prime},t)|\,dt\leq w\,e^{H(M)}\mathsf{L}_{b}(M)\,\|\eta-\eta^{\prime}\|,$
$|\log h_{i}^{\star}(\eta)-\log h_{i}^{\star}(\eta^{\prime})|$ $\leq\frac{|h_{i}^{\star}(\eta)-h_{i}^{\star}(\eta^{\prime})|}{\min h_{i}^{\star}}\leq e^{2H(M)}\mathsf{L}_{b}(M)\,\|\eta-\eta^{\prime}\|.$
We have $2Me^{-H(M)}\leq S(\eta)\leq 2Me^{H(M)}$ and $\mathrm{Lip}_{\eta}(S)\leq 2Me^{H(M)}\mathsf{L}_{b}(M)$, hence
$|a(\eta)-a(\eta^{\prime})|=|\log S(\eta)-\log S(\eta^{\prime})|\leq e^{2H(M)}\mathsf{L}_{b}(M)\,\|\eta-\eta^{\prime}\|.$
@@
$|b_{i}^{\star}(\eta)-b_{i}^{\star}(\eta^{\prime})|\leq 2\,e^{2H(M)}\mathsf{L}_{b}(M)\,\|\eta-\eta^{\prime}\|\quad\Rightarrow\quad\mathrm{Lip}_{\eta}\big(b^{\star}\big)\leq 2\,e^{2H(M)}\mathsf{L}_{b}(M).$
Set $F(\eta)=\big(b^{\star}(\eta),\,v^{\star}(\eta),\,c^{\star}(\eta)\big)\in\mathbb{R}^{4K_{\star}+2},\qquad v_{i}^{\star}(\eta)=u(\eta,z_{i}),\quad c^{\star}(\eta)=T_{p}(\eta,-M).$
So, $b^{\star}$ is $2e^{2H(M)}\mathsf{L}_{b}(M)$-Lipschitz and by Lemma 1 $v^{\star}$ is $\mathsf{L}_{b}(M)$-Lipschitz. We now establish the Lipschitz constant for $c^{\star}$. We start from the quantile identity
@@
By Assumption 1(ii)(ii), for $|\theta|\leq G(M)$ we have $\|\nabla_{\eta}\log p(\theta\mid\eta)\|\leq L_{CP}(G(M))$. Quantile growth ensures $|T_{p}(\eta,z)|\leq G(|z|)\leq G(M)$ for all $z\in(-\infty,-M]$. Therefore
@@
From Assumption 2(ii)(ii), $|\log\partial_{z}T_{p}(\eta,z)|\leq H(|z|)$, hence $\partial_{z}T_{p}(\eta,-M)\leq e^{H(M)}$, so
@@
Combining and applying Mill’s ratio, $c^{\star}$ is $L_{\mathrm{edge}}(M)$-Lipschitz with
@@
By the ReLU vector sup-norm approximation theorem (Yarotsky,, 2017), applied to $F=(b^{\star},v^{\star},c^{\star})$ with
$\mathrm{Lip}_{\eta}(F)\ \leq\ 2e^{2H(M)}\mathsf{L}_{b}(M)+L_{\mathrm{edge}}(M)\qquad\text{where}\qquad L_{\mathrm{edge}}(M)=\frac{e^{H(M)}}{M}\,L_{CP}\!\big(G(M)\big),$
there exists a conditioner $\psi_{\vartheta}:K_{\eta}\mapsto\mathbb{R}^{4K_{\star}+2}$ such that
@@
$m\ \leq\ C_{r}\,(4K_{\star}+2)\,A^{d_{\eta}}\log A,\qquad A\ :=\ \frac{\big(2e^{2H(M)}\mathsf{L}_{b}(M)+L_{\mathrm{edge}}(M)\big)\,\mathrm{diam}(K_{\eta})}{\varepsilon_{\mathrm{par}}}.$
@@
$\delta\coloneqq\sup_{(z,\eta)\in K_{M}}\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|.$
@@
$\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|\leq\mathsf{L}_{b}(M)\,w+\tau,\qquad w=\frac{M}{K_{\star}},$
@@
$K_{\star}=\Bigl\lceil\frac{\mathsf{L}_{b}(M)M}{\varepsilon_{d}}\Bigr\rceil\qquad\text{with}\qquad\varepsilon_{d}=\frac{\varepsilon}{4\,M\,e^{H(M)}},$
so that $\mathsf{L}_{b}(M)w\leq\varepsilon_{d}$.
Then
$\mathsf{L}_{b}(M)w+\tau\ \leq\ \varepsilon_{d}+\varepsilon_{\mathrm{par}}\ \leq\ \frac{\varepsilon}{4Me^{H(M)}}+\frac{\varepsilon}{4Me^{H(M)}}\ =\ \frac{\varepsilon}{2Me^{H(M)}},$
@@
$\sup_{(z,\eta)\in K_{M}}\bigl|\log\partial_{z}T_{\vartheta}(\eta,z)-\log\partial_{z}T_{p}(\eta,z)\bigr|\ \leq\ \varepsilon.$
@@
Using $\delta\leq\varepsilon/(2Me^{H(M)})$ and $e^{x}-1\leq(e-1)x$ for $x\leq 1$ gives $Me^{H(M)}(e^{\delta}-1)\leq\varepsilon(e-1)/2$, and since $\varepsilon_{\mathrm{par}}=\frac{\varepsilon}{4Me^{H(M)}}\leq\varepsilon(e-1)/2$, we have
@@
Thus both inequalities in Assumption 2(iii)(iii) hold with tolerance $\varepsilon$.